(2) Company CDE is financed entirely with equity. The company is thinking of borrowing $900,000. The...

  1. (2) Company CDE is financed entirely with equity. The company is thinking of borrowing $900,000. The loan will be repaid in equal instalments over the next two years, and has a 8% interest rate. The tax rate is 35%. According to the MM proposition with taxes, what would be the increase in the value to CDE after the loan? Show your work.
Answer #1

As per MM proposition with taxes, the increase in value will be the present value of tax shielding of interest of loan.

Tax shielding of Interst Amount of Year 1 = 900,000 * 8% * 0.35 = 25200

Tax shielding of Interst Amount of Year 2 = (900,000 - 450,000) * 8% * 0.35= 12600

***Since half the loan will be paid at the year end 1 450,000 has been subtracted above.

Present Value of Tax Shielding = 25200 / (1+0.08)^1 + 12600/ (1+0.08)^2 = 34,135.80

Increase in value will be $34,135.80